Outline of Nightime Operations (UT Oct 28-30 and Nov 1-2)

Updated to highlight, in red, the tasks which remain to be done and are priorities for any E time that may open up on Wednesday and Thursday, Nov 1 and 2.

If seeing is <~1.2":
  1. LUCI: Verification of AGW transforms.
    1. N30 Field Stop: LUCI1 cannot be configured for N30 Field Stop mask. Configure LUCI1 with N30 Field Stop.
    2. LEFTZEROPOINT: Must update LFBG rotator zeropoint for LUCI1 before verification. LEFTZEROPOINT was -187.917. New LEFTZEROPOINT is -187.315.
    3. Binocular transform collect, see ComTools#Guide_Probe_Transformation for instructions.
      1. Send sync binocular preset ComTools#Sync_Preset
      2. Run transform_collect in two different IDL sessions, one for SX and another for DX (asych presets but to same target so OK).
    4. Use Stone_D_new_Oct2018.list (in /home/LBTO/idl/wfsc/transform/lists)
  2. MODS: Collect data for MODS2 transform.
    1. Run ~modseng/modsScripts/Support/transformInit_bluePoint.pro
      1. Using PointingMask in both MODS.
      2. Use blue channel on MODS2 (and MODS1, for uniformity).
    2. Must update RDG rotator zeropoint before collecting data.
    3. Binocular transform collect.
If seeing is poor:
  1. Collect data for pointing model using MODS/MODS (pointing mask or no pointing mask to get the full 6' square field and put WCS on images?)
  2. Binocular mixed-mode tests, using LUCI1/MODS1 + LBCR
  3. LBC NS guiding test (IT 7387). PCS patch that should have fixed this was backed out on 20181101, though.

LUCI

  1. DONE On-sky checkout of AGW-1 (1 hour)
    • confirm light on lenslets
    • measure/update the rotator center
  2. DONE LUCI-1 measure rotator zeropoint with new astrometric masks (0.5 hour)
  3. DONE LUCI-1 AGW Transform (update the rotator zeropoint for LUCI-2 and can perform LUCI-2 AGW Transform in parallel if mask exchange is done in time (1 hour) - N30 pointing mask (use old one)
  4. ESM testing? Both sides (1 hour)
  5. Verification of AGW Transforms

MODS

  1. Update rotator zero-point for MODS-2 (MODS-1 should be OK)
  2. AGW Transform - requires pointing mask (1 hour)
    1. Stone_P1 and Stone_D fields cover the 1st and 2nd halves of the night. ~/Service_Observing/E_MODSTransforms/. Stone_D PA=270 was done before and Stone_P1 has 3 lists for PA=315.
  3. Verification of AGW Transforms

LBC

  1. Test non-sidereal guiding patch (IT7387)

Telescope Infrastructure/Support

  • DONE Test PCS. GCS, ECS, PSF, OSS patch (various small changes)
  • Apply binodal correction to SX Gregorian (waiting for Doug to analyze the data)
  • collect pointing data with MODS/MODS (~2.5 hours, any seeing)
  • TMS DX M1 Tilt test (10 min)
  • DX M3 alignment at other focal stations
    • for PFU (high)
    • for L-N (high)
    • DONE for LBTI (high)
  • Temperature-free collimation model (from Yang's analysis - not ready as of 10/23)

Binocular Tests (Facility Instruments)

  • Test binocular and pseudomonocular LBC+LUCI/MODS observations:
    • Can we send binocular LBC + LUCI presets via scripts (Olga, a bad-seeing, bad-weather task)
    • Can we send binocular LBC + MODS binocular presets via scripts (Olga, a bad-seeing, bad-weather task)
    • As far as pseudomonocular mode, can IT 5715 be closed? (we believe both are fixed in version 2.15, which is the current version of the LBC software.)
      • when dec=-90, is an initial, non-zero, dither made, or is it ignored?
      • do LBC observations always wait for the optics move to complete?
  • More pointing data after the initial models are tested (MODS first, as LUCI data was taken during restart)
  • DONE Test new version of GCS

AO Tests

  1. Test AO and ESM observations with LUCI-2 to compare with issues found as part of IT 7458 (DX AdSec frequent RIPs of unknown cause)

D-Time Science

  • LMIRCAM/NOMIC (monocular) observations of 29P/Schwassmann-Wachmann (2-3 hours w/overheads) - first part of Sunday
  • MODS LS spectroscopy of stellar merger candidate (2 hours) - first part of Monday
  • MODS LS spectroscopy of nova (1 hour) - end of night (any)
  • MODS LS spectroscopy of Triple AGN candidate (45 min-1hr) - end of night (any)
-- BarryRothberg - 26 Oct 2018
Topic revision: r7 - 01 Nov 2018, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback